# Physics-Constrained Learning of Energy-Preserving Stencils for Maxwell’s Equations

### v3

@@ 6.8 Solver convergence and runtime
Finally, we compare the convergence histories of the different solvers. Figure 9(a) plots the objective value versus iteration for each first-order method for a fixed radius. Figure 9(b) shows the same objective versus wall-clock time. These plots illustrate the tradeoffs among PG, NAG, and ADMM in terms of iterations, per-iteration cost, and constraint enforcement.
(a) Objective vs iteration
(b) Objective vs runtime
@@ 8 Discussion and extensions
To probe more realistic regimes, we also considered (i) a nonstandard target operator whose stencil does not match any classical centered-difference formula and showed that the constrained learner accurately recovers its space–time behavior, while the matching CD stencil incurs noticeable error, and (ii) a noisy regression experiment in which unconstrained least squares produces large coefficients and severe energy drift, whereas the energy- and box-constrained stencil stays well-conditioned, energy-stable, and faithful to the clean baseline $E(x,t)$.
These examples suggest that physics-motivated constraints can regularize noisy or mildly mis-specified regression problems and make it feasible to learn effective Maxwell operators in settings where analytic stencils are difficult to derive or tune (e.g., heterogeneous media or band-limited/experimental data). A natural next step is to extend the same constrained-learning philosophy to staggered curl operators on Yee grids in 2D/3D, enforcing block skew-symmetry to preserve a discrete Poynting theorem, and to couple the learned operators with simple dispersive material models and absorbing layers in more realistic FDTD setups.

### v4

@@ 6.8 Solver convergence and runtime
Finally, we compare the convergence histories of the different solvers. Figure 9 plots the objective value versus iteration and wall-clock time for each first-order method for a fixed radius. These plots illustrate the tradeoffs among PG, NAG, and ADMM in terms of iterations, per-iteration cost, and constraint enforcement.
(a) Objective vs. iteration.
(b) Objective vs. runtime.
@@ 8 Discussion and extensions
To probe more realistic regimes, we also considered a nonstandard target operator whose stencil does not match any classical centered-difference formula and showed that the constrained learner accurately recovers its space–time behavior, while the matching CD stencil incurs noticeable error, and a noisy regression experiment in which unconstrained least squares produces large coefficients and severe energy drift, whereas the energy- and box-constrained stencil stays well-conditioned, energy-stable, and faithful to the clean baseline $E(x,t)$.
These examples suggest that physics-motivated constraints can regularize noisy or mildly mis-specified regression problems and make it feasible to learn effective Maxwell operators in settings where analytic stencils are difficult to derive or tune (e.g., heterogeneous media or band-limited/experimental data). A natural next step is to extend the same constrained-learning philosophy to staggered curl operators on Yee grids in higher dimensions, enforcing block skew-symmetry to preserve a discrete Poynting theorem, and to couple the learned operators with simple dispersive material models and absorbing layers in more realistic FDTD setups.
@@ Acknowledgements
Appendix: Spectral structure of the learned stencil
Theorem 3 (Spectral structure and modal energy conservation for the learned stencil).
Let $w^{\star}\in\mathbb{R}^{2R+1}$ be any solution of the quadratic program (19) (or the equality-constrained variant (20)), and let $D^{\star}$ be the associated periodic convolution operator defined by (2). Let $\mu^{\star}(\theta)$ denote the Fourier symbol of $D^{\star}$,
$\mu^{\star}(\theta):=\sum_{k=-R}^{R}w_{k}^{\star}\,\mathrm{e}^{\mathrm{i}k\theta},\qquad\theta\in[-\pi,\pi].$
Then the following hold:
  1. (i)
    Because $w^{\star}$ satisfies the linear constraint $Cw^{\star}=d$, it is skew-symmetric in the sense of (10), i.e.
    $w_{0}^{\star}=0,\qquad w_{-\ell}^{\star}=-w_{+\ell}^{\star},\quad\ell=1,\dots,R.$
    Consequently, the symbol $\mu^{\star}(\theta)$ is purely imaginary and odd:
    $\mu^{\star}(\theta)\in\mathrm{i}\mathbb{R},\qquad\mu^{\star}(-\theta)=-\mu^{\star}(\theta)\quad\text{for all }\theta\in[-\pi,\pi].$
  2. (ii)
    Consider the semi-discrete Maxwell system with the learned operator $D^{\star}$,
    $\frac{d}{dt}\begin{bmatrix}\mathbf{H}\\[2.5pt] \mathbf{E}\end{bmatrix}=A^{\star}\begin{bmatrix}\mathbf{H}\\[2.5pt] \mathbf{E}\end{bmatrix},\qquad A^{\star}=\begin{bmatrix}0&D^{\star}\\ (D^{\star})^{\top}&0\end{bmatrix}$,
    together with the discrete $L^{2}$ inner product (4) and energy (5). Then $A^{\star}$ is skew-adjoint and diagonalizable in the discrete Fourier basis. For each discrete wavenumber $\theta$, the corresponding $2\times 2$ block has eigenvalues
    $\lambda_{\pm}^{\star}(\theta)=\pm\,\mathrm{i}\,|\mu^{\star}(\theta)|.$
    In particular, each Fourier mode evolves as a harmonic oscillator with frequency $|\mu^{\star}(\theta)|$.
  3. (iii)
    The discrete electromagnetic energy associated with the learned stencil can be written as a sum of modal contributions and is conserved mode by mode. If $\widehat{E}^{\star}(\theta,t)$ and $\widehat{H}^{\star}(\theta,t)$ denote the discrete Fourier coefficients of $\mathbf{E}(t)$ and $\mathbf{H}(t)$, then
    $\mathcal{E}(t)=\frac{\Delta x}{2}\sum_{\theta}\bigl(|\widehat{E}^{\star}(\theta,t)|^{2}+|\widehat{H}^{\star}(\theta,t)|^{2}\bigr),$
    and for each $\theta$ one has
    $\frac{d}{dt}\bigl(|\widehat{E}^{\star}(\theta,t)|^{2}+|\widehat{H}^{\star}(\theta,t)|^{2}\bigr)=0.$
    Thus the total discrete energy and each modal energy contribution are preserved exactly by the semi-discrete learned Maxwell system.
Proof.
Because $w^{\star}$ is a solution of (19) or (20), it satisfies the linear constraint $Cw^{\star}=d$ by construction. By Corollary 1, this is equivalent to the skew-symmetry conditions (10), which immediately imply the structure of $\mu^{\star}(\theta)$ and the identities in part (i) via the same calculation as in Theorem 2.
Skew-symmetry of $w^{\star}$ implies $(D^{\star})^{\top}=-D^{\star}$, so the block matrix $A^{\star}$ is skew-adjoint with respect to the discrete inner product (4). Since $D^{\star}$ is a periodic convolution, it is diagonalized by the discrete Fourier transform; the $2\times 2$ block for each mode $\theta$ has the form
$A^{\star}(\theta)=\begin{bmatrix}0&\mu^{\star}(\theta)\\ -\mu^{\star}(\theta)&0\end{bmatrix}$,
whose eigenvalues are $\lambda_{\pm}^{\star}(\theta)=\pm\,\mathrm{i}\,|\mu^{\star}(\theta)|$, yielding part (ii).
Finally, unitarity of the discrete Fourier transform implies that the discrete $L^{2}$ norms of $\mathbf{E}$ and $\mathbf{H}$ are equal to the sums of squared modal amplitudes, so the energy $\mathcal{E}(t)$ decomposes as in part (iii). For each fixed $\theta$, the $2\times 2$ system governed by the skew-Hermitian matrix $A^{\star}(\theta)$ preserves the Euclidean norm of $(\widehat{H}^{\star}(\theta,t),\widehat{E}^{\star}(\theta,t))$, which implies modal energy conservation and hence conservation of the total discrete energy. ∎